OpenStudy (anonymous):

My last question- this one I'm not too sure of. --------------------------------------

OpenStudy (anonymous):

heptagon has 7 sides you have to find the total sum of angles in the shape 180(n-2) = 180 * 5 = 900 subtract 144 from 900 then 6x = 756 x= 126

jimthompson5910 (jim_thompson5910):

the steps and final answer are correct nice work

OpenStudy (anonymous):

Thanks for checking :)

jimthompson5910 (jim_thompson5910):

np